OpenStudy (anonymous):

Simplify each radical expression. Use absolute value symbols as needed. √400x^2y^6

OpenStudy (anonymous):

\[\sqrt{400x^2y^6}=\sqrt{400}\sqrt{x^2}\sqrt{y^6}=20|xy^3|\]